OpenStudy (anonymous):

Use a metric ruler and a protractor to find 2a--> - 2b-->, if |a-->| = 2cm and |b-->| = 1cm. Then find the magnitude and direction of the resultant.

OpenStudy (anonymous):

I don't really get how I would be doing this...

OpenStudy (anonymous):

Possible answers: 8 cm, 41° 9 cm, 38° 6 cm, 36° 4 cm, 31°

OpenStudy (anonymous):

@jim_thompson5910 @whpalmer4

jimthompson5910 (jim_thompson5910):

I think you would use the idea that if vector a = <x,y> then x = r*cos(theta) x = |a|*cos(60) and y = r*sin(theta) y = |a|*sin(60)

OpenStudy (anonymous):

Thank you c:

jimthompson5910 (jim_thompson5910):

you're welcome tell me what you get

OpenStudy (anonymous):

Alrighty c:

OpenStudy (anonymous):

What is r btw?

jimthompson5910 (jim_thompson5910):

r is the length of the vector so r = |a| or r = |b|

OpenStudy (anonymous):

ah okai c:

jimthompson5910 (jim_thompson5910):

basically if you drew a circle around the tip of the vector, it would be the radius of that circle (the vector is centered at (0,0))

OpenStudy (anonymous):

Then for the y part did you mean |b| or |a|?

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

no it applies to both

jimthompson5910 (jim_thompson5910):

r isn't the same for both vectors, it's just in general

jimthompson5910 (jim_thompson5910):

in the case of vector a, r = 2 in the case of vector b, r = 1

OpenStudy (anonymous):

Oh okai so when I do both those x equations, do I add them together or do I multiply them?

jimthompson5910 (jim_thompson5910):

x = r*cos(theta) x = |a|*cos(60) x = 2*cos(60) x = 2*(1/2) x = 1 y = r*sin(theta) y = |a|*sin(60) y = 2*sin(60) y = 2*(sqrt(3)/2) y = sqrt(3) So the vector a has the terminal point at ( 1, sqrt(3) )

OpenStudy (anonymous):

So cos(60) = 1/2?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

For vector b... x = r*cos(theta) x = |b|*cos(135) x = 1*cos(135) x = 1*(-sqrt(2)/2) x = -sqrt(2)/2 y = r*sin(theta) y = |b|*sin(135) y = 1*sin(135) y = 1*(sqrt(2)/2) y = sqrt(2)/2 and the vector b has the terminal point at ( -sqrt(2)/2 , sqrt(2)/2 )

OpenStudy (anonymous):

Oh damn, I was using radians >_<

jimthompson5910 (jim_thompson5910):

ah that gets me everytime

jimthompson5910 (jim_thompson5910):

once you've found the x,y components, you can add and subtract the vectors 2a - 2b = 2(a-b) x coordinate of vector a-b: 1 - (-sqrt(2)/2) = 1+sqrt(2)/2 y coordinate of vector a-b: sqrt(3) + sqrt(2)/2 = 1+sqrt(2)/2 Now double everything to find the components for the vector 2(a-b) x coordinate of vector 2(a-b): 2(1+sqrt(2)/2) = 2+sqrt(2) y coordinate of vector 2(a-b): 2(sqrt(3) + sqrt(2)/2) = 2*sqrt(3)+sqrt(2)

jimthompson5910 (jim_thompson5910):

So effectively, we know that x = 2+sqrt(2) y = 2*sqrt(3)+sqrt(2) the last thing to do is plug them into the formulas r = sqrt(x^2 + y^2) and theta = arctan(y/x) to find the magnitude and the direction

OpenStudy (anonymous):

Okai 1 quick question before I read what you said, you know how you got 2 for 2*cos theta and 1 for 2*cos(60)? Did you minus them to get the final 1 for the x?

jimthompson5910 (jim_thompson5910):

yes I did these basic steps found the x,y component form for vector a found the x,y component form for vector b once you have the component form, you can easily add/subtract the vectors the resultant vector a-b is made up of x,y components as well and the x component is found by subtracting the corresponding x components for vectors 'a' and b (y is found the same way)

OpenStudy (anonymous):

Okai good to know, I want to know I can do it later on with another problem that's why I was asking c:

jimthompson5910 (jim_thompson5910):

oops made a typo y coordinate of vector 2(a-b) *should* be 2*sqrt(3)-sqrt(2)

jimthompson5910 (jim_thompson5910):

but everything else looks ok

OpenStudy (anonymous):

Okai so for r=sqrt (x^2+y^2) I got 4.06 which I rounded to 4cm. Okai so now I know my answer, but how would I get 31degrees?

jimthompson5910 (jim_thompson5910):

use theta = arctan(y/x)

OpenStudy (anonymous):

Arctan is just [] right? Or am I wrong?

jimthompson5910 (jim_thompson5910):

what do you mean just []

OpenStudy (anonymous):

Oh nevermind it's inverse tan right?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

arctangent = inverse of the tangent function

OpenStudy (anonymous):

ohhh sorry i didn't know

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (anonymous):

I'm not getting the right answer for that part :l I've got 22.9, 55.01, and 1.06 *me trying it different ways*

jimthompson5910 (jim_thompson5910):

theta = arctan(y/x) theta = arctan( (2*sqrt(3)-sqrt(2))/(2+sqrt(2))) theta = 30.9805340840034 theta = 31 (rounding to the nearest degree)

OpenStudy (anonymous):

oh i forgot the extra ( in the front

jimthompson5910 (jim_thompson5910):

hmm odd how there wasn't an error

OpenStudy (anonymous):

ohh i used + as well instead of -

jimthompson5910 (jim_thompson5910):

that's my bad lol i had + but i edited it to -

jimthompson5910 (jim_thompson5910):

not sure how i got + when a-b tells us we're subtracting components

OpenStudy (anonymous):

lmao so the first answer is wrong? cx

jimthompson5910 (jim_thompson5910):

no, the first answer of 4 is correct

jimthompson5910 (jim_thompson5910):

notice how the 4 cm and the 31 degrees match up and it's choice D

OpenStudy (anonymous):

oh so the first one i add in the y but the second one i subtract in the y?

jimthompson5910 (jim_thompson5910):

no it's all supposed to be subtracting (in the numerator)

jimthompson5910 (jim_thompson5910):

before, i mistakenly got this y = 2*sqrt(3)+sqrt(2) when it should have been this y = 2*sqrt(3)-sqrt(2)

OpenStudy (anonymous):

But for the first one I used the + and got the correct answer of 4cm

jimthompson5910 (jim_thompson5910):

you got close to the correct answer, but rounding made you get there anyway

OpenStudy (anonymous):

Well yeah but you know what I mean cx

jimthompson5910 (jim_thompson5910):

I got 3.98232284958912 for the magnitude which is much closer that rounds to 4 as well

OpenStudy (anonymous):

Oh okai

OpenStudy (anonymous):

I get you c:

jimthompson5910 (jim_thompson5910):

ok that's great

OpenStudy (anonymous):

Thank you n.n

jimthompson5910 (jim_thompson5910):

np